I need help questions the writes
express as sum,difference and multiples of logarithms and simplify
log 3 <--lowered(3x^2)

Answer by stanbon(75887)

log3(3x^2)
-------
= log3(3) + log3(x^2)
-------
= 1 + 2log3(x)
-------
2log3(x) = -1
log3(x) = -1/2
x = 3^(-1/2)
---
x = (1/3)^(1/2)
---
x = sqrt(1/3)
================
Cheers,
Stan H.
================